At a community college, the difference between a professor's salary and an associate professor's salary is $13,000. The sum of their salaries is $99,000. Find the salary of an associate professor at the college.
Answer by stanbon(75887)

At a community college, the difference between a professor's salary and an associate professor's salary is $13,000. The sum of their salaries is $99,000. Find the salary of an associate professor at the college.
-----
Equations:
p - a = 13000
p + a = 99,000
-----------
Add and solve for "p"::
2p = 112,000
p = $56,000 (professor's salary)
---
Solve for "a":
a = 99000-56000
a = $43,000
------------------
Cheers,
Stan H.
-----------